OpenStudy (anonymous):

Kiera wants to increase her savings account to a balance of $1,350. Her current balance is $174 and she adds $84 a month to her savings. How many months will it take for her to reach her goal of $1,350? 16 months 14 months 18 months 12 months

OpenStudy (igreen):

We can use the equation: 174 + 84x = 1350. Plug in 16 months: 174 + 84(16) = 1350 Multiply: 174 + 1344 = 1350 1518 = 1350 16 months is incorrect. Plug in 14 months: 174 + 84x = 1350 174 + 84(14) = 1350 174 + 1176 = 1350 1350 = 1350 So 14 months is correct!

OpenStudy (anonymous):

174+(84x14)=1,350 is your answer make the answer reasonable by adding and multiplying in one equation